Please help explain this!

Please help explain this!
Answer #1

V = kq/r

q = V*r/k

A = 4*pi*r^2

r = sqrt(A/(4*pi))

= sqrt (1.28/(4*pi)

= 0.32 m

q = 587*0.32/(9*10^9)

= 2.09*10^-8 C